OpenStudy (anonymous):

help me sqrt(5+2sqrt6)+sqrt(8-2sqrt5)

OpenStudy (anonymous):

simplify it

OpenStudy (anonymous):

\[\LARGE \sqrt{5+2\sqrt6}+\sqrt{8-2\sqrt5}\]

OpenStudy (anonymous):

yes this is the problem

OpenStudy (anonymous):

I can't simplify it any further...

OpenStudy (experimentx):

Try rationalizing it.

OpenStudy (anonymous):

i tried but how experiment x can u give me some hint

OpenStudy (anonymous):

i don't think you can do a thing with the second term

OpenStudy (experimentx):

Oh ...

OpenStudy (anonymous):

the first one is not too bad since \[5+2\sqrt{6}=2+2\sqrt{6}+3=(\sqrt{2}+\sqrt{3})^2\] meaning \[\sqrt{5+2\sqrt{6}}=\sqrt{2}+\sqrt{3}\] maybe there is a similar trick for the second one i don't see it yet

OpenStudy (anonymous):

here is the sol. that i got from bagatrix

OpenStudy (anonymous):

can any1 explain me this ?

OpenStudy (anonymous):

something goofy in this solution.

OpenStudy (anonymous):

@anhkhoavo1210 i don't think so

OpenStudy (anonymous):

sr i deleted it :D wait for 5 minutes

OpenStudy (anonymous):

what is <R><r>? and i assume there is typesetting problem because i one line i see \(\sqrt{8-25}\) which must be something else, maybe \(\sqrt{8-2\sqrt{5}}\)

OpenStudy (anonymous):

yes i am thinking the same for hours is this a wrong question

OpenStudy (anonymous):

i think it must be \[8 - 2\sqrt 15\] :)

OpenStudy (anonymous):

\[\sqrt {15}\]

OpenStudy (anonymous):

i can't get it. i got \[(\sqrt{5}-1)^2=6-2\sqrt{5}\] but i am not getting it with the 8

OpenStudy (anonymous):

could it be with fractions something? o.O

OpenStudy (anonymous):

ok first of all that solution is wrong, you can check by a calculator

OpenStudy (anonymous):

or by experimentx link

OpenStudy (anonymous):

thanks for proving it false experiment atleast 1 soln is false. now what is the true result ?

OpenStudy (anonymous):

the given MUST be wrong... (I think..)

OpenStudy (anonymous):

so forget that solution typos and all i cannot see what to do with this

OpenStudy (mertsj):

The only thing I can see is what satellite said for the first term and then for the second one: \[\sqrt{8-2\sqrt{5}}=\sqrt{2(4-\sqrt{5)}}=\sqrt{2}\sqrt{4-\sqrt{5}}\] but what good is that? None

OpenStudy (phi):

here is the sol. that i got from bagatrix If you evaluate the answer sqrt(7)+sqrt(6) you see it does not match the original expression.

OpenStudy (angela210793):

I don't think so either....:P i think i might have multiplied the numbers in bracket forgetting the squre roots :S